# revtex4-2Repair the float

### v1

@@ Abstract
We propose a new approach to simulate neutrino scattering events as an alternative to the standard Monte Carlo generator approach. Generative adversarial neural network (GAN) models are developed to simulate neutrino-carbon collisions in the few-GeV energy range. The models produce scattering events for a given neutrino energy. GAN models are trained on simulation data from NuWro Monte Carlo event generator. Two GAN models have been obtained: one simulating only quasielastic neutrino-nucleus scatterings and another simulating all interactions at given neutrino energy. The performance of both models has been assessed using two statistical metrics. It is shown that both GAN models successfully reproduce the event distributions.
@@ I Introduction
Thus, oscillation analyses require procedures to reconstruct neutrino energy from the detected final-state particles. These procedures rely heavily on Monte Carlo (MC) simulations [5], which must incorporate models for all relevant physical processes, including various types of interactions on bound nucleons (or, at higher energies, on quarks), contributions from two-body currents, and coherent neutrino-nucleus reactions. Additionally, final state interactions (FSI), i.e. secondary interactions of hadrons inside nuclei, must be accounted for.
@@
MC generators for events (in particle and nuclear physics) serve as a bridge between theory and experiment [11]. Monte Carlo simulations are employed at every stage of an experimental study, including the production of the neutrino flux, interactions within a detector, and the detector’s response. In this paper, we focus entirely on neutrino interactions inside the detector.
The MC generators, equipped with implementations of theoretical and phenomenological models, are often tuned to match experimental data. Theoretical models implemented in generators typically rely on uncertain internal parameters adjusted using available experimental data. However, these procedures cannot fully account for limitations stemming from theoretical models, which involve many approximations, particularly in the case of nuclear effects. Therefore, it would be beneficial to develop alternatives to existing MC event generators that can learn directly from available experimental cross-section data and modify their predictions accordingly. As a result, efforts have been made to explore new approaches for developing alternatives to standard MC generators. A novel approach has recently emerged: leveraging advanced artificial intelligence (AI) models to dynamically adapt to experimental data, enabling highly effective simulations of physical phenomena. Machine-learning methods have been applied to study and model neutrino-nucleus scattering and related electron-nucleus processes. These applications include extracting information about the axial content of the nucleon [12], modeling electron-nucleus scattering cross-sections [13, 14, 15, 16], and accelerating sampling from given differential cross-section distributions [17].
This paper explores, adapts, and advances modern machine-learning techniques for simulating neutrino-nucleus scattering events. We focus on generative adversarial networks (GANs) [18]. The GAN technique enables training a neural network that learns to generate data produced by an unknown or poorly understood mechanism. A successful generative model takes an input latent vector drawn from a given distribution in the latent space (typically a Gaussian distribution) and generates samples that match the characteristics of the training data.
@@
This paper applies the GAN technique to simulate charged current (CC) neutrino-carbon scattering. We consider two cases. First, we test the applicability of our approach to quasielastic (QE) scattering. Then, we extend our study to include all interaction modes at a given neutrino energy. In both cases, the GAN takes neutrino energy as input and generates the charged lepton kinematics. The model is applicable to neutrino energies ranging from a few hundred MeV to tens of GeV, which are typical for neutrino oscillation experiments.
@@ II Method
We aim to obtain the deep neural network (DNN) that generates neutrino-nucleus scattering events. As an illustrative example, we consider muon neutrino charged current interactions with a carbon target. The DNN model is trained (optimized) to learn the underlying probability distribution. As a source of information on neutrino-carbon scattering, we employ the NuWro MC generator.
@@
The procedure for getting the model is the following:
@@
II.1 Data and kinematic settings
We consider CC QE and inclusive (INC) scattering of muon neutrinos off carbon. The QE and INC data are obtained from a simulation done with the NuWro MC event generator, using the spectral function approach to model the nucleus in the QE channel [34] and the local Fermi gas in other channels. In the QE case, the generated events contain muons and nucleons or sometimes pions due to the final state interaction effects. In the INC case, the final states contain muons and a variety of mesons and baryons. In both types of events, this study only considers the muon’s kinematics.
Let us denote the neutrino energy by $E_{\nu}$. The outgoing muon is characterized by its energy $E_{\mu}$ and momentum $\mathbf{p}_{\mu}$. The angle between the neutrino and muon momenta is denoted by $\theta$.
@@ II.1 Data and kinematic settings
The neural networks work most efficiently when input variables are in the range $\sim(-1,1)$. Hence, we introduce the re-scaled energy variable:
@@
As explained above, the network generates two variables that define the muon kinematics. After considering various options, we use a pair $(E_{\mu}^{\prime},\theta^{\prime})$, given by:
@@
where $E^{-}_{\nu}=E_{\nu}-9.8$ MeV is used to ensure a tighter upper bound. Both variables take values ranging approximately from $-1$ to $1$.
@@ II.2 GAN for event generation
We also employ a variation of Block 1 (2), called Block 3 (4), in which the information about the condition $E^{\prime}_{\nu}$ is concatenated to its input right before the dense layer, as shown in Fig. 3. Block 1(2) with extra $E^{\prime}_{\nu}$ input will be called Block 3(4), respectively. Note that these blocks are used only by INC models.
@@
The input of D-QE(INC) is either the NuWro sample vector $(\mathbf{v},E^{\prime}_{\nu})$ or the output of the generator $(G(\mathbf{x},E^{\prime}_{\nu}),E^{\prime}_{\nu})$. The input is transformed in parallel by two branches of the same architecture (see Fig. 4(b)). Each branch consists of a sequence of two Blocks 1. Then, the outputs of both branches are concatenated and transformed by the main body, which is made by one Block 1(3) and a sequence of three Blocks 2(4). The neural network has a total depth of six Blocks. The last section of the discriminator is a fully connected layer with a single neuron and linear activation function (a sigmoid is applied on the output when it contributes to the loss).
@@
The blocks of the G-QE and the D-QE have $100$ neurons, except for the first Block 1 of G-QE, which has $200$ neurons, and the two dense layers of each initial pair of branches of the D-QE, which consists of $50$ neurons. Similarly, the G-INC comprises blocks and dense layers of $100$ neurons, but the first block has $200$ neurons (just as the G-QE). The D-INC model has more neurons in the blocks than its QE counterpart. In D-INC, most blocks consist of $141$ neurons, but the first two dense layers in the initial branches have $70$ neurons each.
@@ II.3 Optimization scheme
$L_{D,1}(G)$ $=$ $-\frac{1}{B}\sum_{i=1}^{B}\log[D(G(\mathbf{x}_{i},E_{\nu,i}),E_{\nu,i}]$ (5)
$L_{D,2}(\textbf{v})$ $=$ $-\frac{1}{B}\sum_{j=1}^{B}\log[1-D(\textbf{v}_{j},E_{\nu,j})]$ (6)
$L_{G}(G)$ $=$ $-\frac{1}{B}\sum_{k=1}^{B}\log[1-D(G(\mathbf{x}_{k}),E_{\nu,k})],$ (7)
@@
GANs are optimized using the so-called mini-max framework [43]. Hence, a successful model is a result of a balance between two losses: of the discriminator (5) and (6) and of the generator (7). The Eq. (5) is computed on “fake” data generated by G, whereas the loss (6) is calculated on “true” samples obtained from NuWro. The last loss (7) is evaluated on generated by model “fake” samples.
@@ II.4 Testing model’s quality
Figure 7: Top row: NuWro (left) and GAN (right) generated samples for the QE scattering. Bottom row: Pull w/o tails per bin from NuWro and G-QE samples (left) and pull distribution (right).
@@
where $n_{\text{\sc NuWro}{},i}$ and $n_{gan,i}$ are the bin contents at the $i$-th bin, and $\sigma_{\text{\sc NuWro}{},i}^{2}=n_{\text{\sc NuWro}{},i}$ and $\sigma_{gan,i}^{2}=n_{gan,i}$ are statistical uncertainties. The bin contents of a histogram are expected to follow the Poisson distribution. As the G event distribution gets closer to NuWro’s, then Eq. (8) defines random variable that follows a standard normal distribution $N(0,1)$ for large $n_{\text{\sc NuWro}{},i}$ and $n_{gan,i}$. For a random variable $x\sim N(0,1)$, the expected value of $|x|$ is
@@
The MAP metric is computed using histograms of the joint $E_{\mu}^{\prime}$ and $\theta^{\prime}$ distribution, as shown in Figs. 7 and 8. To reduce the bias coming from extremely low statistics bins at distribution “tails”, we compute the MAP for bins for which $n_{nuw,i}>5$ and $n_{gan,i}>5$.
@@
In Fig. 9, we show how the EMD and MAP metrics change during the training of the QE and INC models. As can be noticed, after $100$ epochs, the optimization leads to the model for which the metrics begin to saturate. We stop the optimization using the MAP and EMD metrics. The best model minimizes both.
@@ III Results
Table 1: EMD and MAP (with and without tails) computed for the samples generated by the NuWro and the G-QE for fixed neutrino energy $E_{\nu}$ and randomly sampled from a uniform $E_{\nu}$ distribution (“All” tag).
@@
[…]
Figure 11: Event histograms, depending on $\theta^{\prime}$ (left column) and $\theta$ (right column) generated by the G-QE and NuWro for neutrino energies =0.5absent0.5=0.5= 0.51111,  and $5$ GeV, shown in top, middle, and bottom rows, respectively.
Figure 12: Event histograms, depending on $p_{\mu,z}$ generated by the G-QE and NuWro for neutrino energies 0.50.50.50.51111, and $3$ GeV, shown in top left, top right, and bottom figures, respectively.
The quality of the QE-GAN model and its dependence on neutrino energy is presented in Table 1. We provide the values of the EMD and MAP metrics computed for several values of neutrino energies, between $0.5$ and $9$ GeV, as well as for a dataset with energies distributed uniformly in the energy region that covers the same energy domain as the training dataset. The performance of the G-QE model is excellent in all the considered cases.
In Fig. 7, we show the two-dimensional histograms and pull distribution of events generated by the NuWro and G-QE that are uniformly distributed in energies from $300$ MeV to $10$ GeV. The pulls per bin resemble pure noise, and their distribution approaches the Gaussian shape.
@@
Figure 13: Event histograms, depending on $E_{\mu}^{\prime}$ (left column) and $E_{\mu}$ (right column) generated by the G-INC and the NuWro for neutrino energies =1absent1=1= 13333, and $5$ GeV, shown in top, middle and bottom rows, respectively.
[…]
Additionally, Fig. 11 presents the dependence on $\theta^{\prime}$ and $\theta$ for neutrino energies of 0.50.50.50.51111, and $5$ GeV. Furthermore, we derive the longitudinal momentum as a composite variable (1) and show its event distribution in Fig. 12. The agreement between the GAN and the NuWro generated results is excellent. Note that for the comparison, we generated a new set of the NuWro events that were not used during training for all these tests.
Let us now discuss the INC model. As for the QE analysis, in Fig. 8, we show the two-dimensional histograms and pull distribution of INC events generated by the NuWro and G-INC that are uniformly distributed in energies from $300$ MeV to $10$ GeV. Again, a good agreement is achieved between the NuWro and G-INC generator. In Fig. 13, we plot the histograms of events generated by the G-INC and the NuWro MC generator for neutrino energies 11113333 and $5$ GeV for both $E^{\prime}_{\mu}$ and $E_{\mu}$.
To test if our GAN model can reproduce the QE and $\Delta(1232)$ resonance peaks equally well, let us now consider event distributions for hadronic invariant mass $W$ defined by
@@
where $\omega=E_{\nu}-E_{\mu}$ is the energy transferred to the nucleus, $M_{N}$ the average mass of the nucleon, and $Q^{2}$ the four-momentum transfer
@@
If one disregards the Fermi motion of nucleons in the target nucleus and their binding energies effect, the value of $W$ can be obtained from the muon kinematics alone.

### v2

# Generative adversarial neural networks for simulating neutrino interactions

@@ Abstract
We propose a new approach to simulate neutrino scattering events as an alternative to the standard Monte Carlo generator approach. Generative adversarial neural network (GAN) models are developed to simulate charged current neutrino-carbon collisions in the few-GeV energy range. We consider a simplified framework to generate muon kinematic variables, specifically its energy and scattering angle. GAN models are trained on simulation data from NuWro Monte Carlo event generator. Two GAN models have been obtained: one simulating quasielastic neutrino-nucleus scatterings and another simulating all interactions at given neutrino energy. The models work for neutrino energy ranging from 300 MeV to 10 GeV. The performance of both models has been assessed using two statistical metrics. It is shown that both GAN models successfully reproduce the distribution of muon kinematics.
@@ I Introduction
Thus, oscillation analyses require procedures to reconstruct neutrino energy from the detected final-state particles. These procedures rely heavily on MC simulations [5], which must incorporate models for all relevant physical processes, including various types of interactions on bound nucleons (or, at higher energies, on quarks), contributions from two-body currents, and coherent neutrino-nucleus reactions. Additionally, final state interactions (FSI), i.e. secondary interactions of hadrons inside nuclei, must be accounted for.
@@
MC generators of events in particle and nuclear physics serve as a bridge between theory and experiment [11]. Monte Carlo simulations are employed at every stage of an experimental study, including the production of the neutrino flux, interactions within a detector, and the detector’s response. In this paper, we focus entirely on neutrino interactions inside the detector.
The MC generators, equipped with implementations of theoretical and phenomenological models, are often tuned to match experimental data. Theoretical models implemented in generators typically rely on uncertain parameters adjusted using available experimental data. However, these procedures cannot fully account for limitations stemming from theoretical models, which involve many approximations, particularly in the case of nuclear effects. Therefore, it would be beneficial to develop alternatives to existing MC event generators that can learn directly from available experimental cross section data and modify their predictions accordingly. As a result, efforts have been made to explore new approaches for developing alternatives to standard MC generators. A novel approach has recently emerged: leveraging advanced artificial intelligence (AI) models to dynamically adapt to experimental data, enabling highly effective simulations of physical phenomena. Machine-learning methods have been applied to study and model neutrino-nucleus scattering and related electron-nucleus processes. These applications include extracting information about the axial content of the nucleon [12], modeling electron-nucleus scattering cross sections [13, 14, 15, 16], and accelerating sampling from given differential cross section distributions [17, 18].
This paper explores, adapts, and advances modern machine-learning techniques for simulating neutrino-nucleus scattering events. We focus on generative adversarial networks (GANs) [19]. The GAN technique enables training a neural network that learns to generate data produced by an unknown or poorly understood mechanism. A successful generative model takes as input a latent vector drawn from a given distribution in the latent space (typically a Gaussian distribution) and generates samples that match the characteristics of the training data.
@@
This paper applies the GAN technique to simulate charged current (CC) neutrino-carbon scattering. We adopt a simplified framework wherein the models are tasked with generating muon kinematic variables. We consider two cases. First, we test the applicability of our approach to quasielastic (QE) scattering. Then, we extend our study to include all interaction modes at a given neutrino energy. In both cases, the GAN takes neutrino energy as input and generates the charged lepton kinematics. The model is applicable to neutrino energies ranging from a few hundred MeV to tens of GeV, which are typical for neutrino oscillation experiments.
@@ II Method
We aim to obtain a deep neural network (DNN) that generates neutrino-nucleus scattering events. As an illustrative example, we consider muon neutrino charged current interactions with a carbon target. The DNN model is trained (optimized) to learn the underlying probability distribution. As a source of information on neutrino-carbon scattering, we employ the NuWro MC generator.
@@
The procedure for developing the model is the following:
@@ II.1 Data and kinematic settings
We consider CC QE and inclusive (INC) scattering of muon neutrinos off carbon. The QE and INC data are obtained from a simulation done with the NuWro MC event generator, using the spectral function approach to model the nucleus in the QE channel [36] and the local Fermi gas in other channels. In the QE case, the generated events contain muon and nucleons or sometimes pions due to the final state interaction effects. In the INC case, the final states contain muons and a variety of mesons and baryons. In both types of events, this study only considers the muon’s kinematics.
Let us denote the neutrino energy by $E_{\nu}$. The outgoing muon is characterized by energy $E_{\mu}$ and momentum $\mathbf{p}_{\mu}$. The angle between the neutrino and muon momenta is denoted by $\theta$.
@@
The neural networks work most efficiently when input variables are in the range $\sim(-1,1)$. Hence, we introduce the rescaled energy variable:
@@
As explained above, the network generates two variables that define the muon kinematics. After considering various options, we use a pair $(E_{\mu}^{\prime},\theta^{\prime})$, given by [^1] We explored various transformations of the kinematic variables to optimize the performance of our neural network model. We aimed to transform the data into a format where the model operates most efficiently and where optimization yields the best results. We chose a range of [-1, 1] because in this domain, typical activation functions perform most effectively. Additionally, we applied a square root transformation to reduce skewness in the event distributions, which aids the training process.:
@@
where $E^{-}_{\nu}=E_{\nu}-9.8$ MeV is used to ensure a tighter upper bound [^2] It effectively takes into account the effect of binding energy.. Both variables take values in the range approximately from $-1$ to $1$.
@@ II.2 GAN for event generation
We also employ a variation of Block 1 (2), called Block 3 (4), in which the information about the condition $E^{\prime}_{\nu}$ is concatenated to its input right before the dense layer, as shown in Fig. 3. Block 1(2) with extra $E^{\prime}_{\nu}$ input will be called Block 3(4), respectively. Note that these blocks are used only by INC models [^4] The inclusive data is more complex than the QE data. Initially, we considered using the same neural network model to generate both INC and QE events. However, we could not get satisfactory results for the INC data, and we ultimately decided to implement some modifications to better capture the more intricate structure of the data..
@@
The input of D-QE(INC) is either the NuWro sample vector $\left(\mathbf{v},E^{\prime}_{\nu}\right)$ or the output of the generator $\left(G(\mathbf{x},E^{\prime}_{\nu}),E^{\prime}_{\nu}\right)$. The input is transformed in parallel by two branches of the same architecture (see Fig. 4(b)). Each branch consists of a sequence of two Block 1 layers. Then, the outputs of both branches are concatenated and processed by the main body, composed of one Block 1(3) and a sequence of three Block 2(4) layers. The neural network has a total depth of six Block layers. The final section of the discriminator is a fully connected layer with a single neuron and a linear activation function (a sigmoid is applied to the output when it contributes to the loss).
@@
The blocks of the G-QE and the D-QE have $100$ neurons, except for the first Block 1 of G-QE, which has $200$ neurons, and two dense layers of each initial pair of branches of the D-QE, which consists of $50$ neurons. Similarly, the G-INC comprises blocks and dense layers of $100$ neurons, but the first block has $200$ neurons (as the G-QE). The D-INC model has more neurons in the blocks than its QE counterpart. In D-INC, most blocks consist of $141$ neurons, but the first two dense layers in the initial branches have $70$ neurons each.
@@ II.3 Optimization scheme
$L_{D,1}(\textbf{v})$ $=$ $-\frac{1}{B}\sum_{j=1}^{B}\log[1-D(\textbf{v}_{j},E_{\nu,j})],$ (5)
$L_{D,2}(G)$ $=$ $-\frac{1}{B}\sum_{i=1}^{B}\log[D(G(\mathbf{x}_{i},E_{\nu,i}),E_{\nu,i})],$ (6)
$L_{G}(G)$ $=$ $-\frac{1}{B}\sum_{k=1}^{B}\log[1-D(G(\mathbf{x}_{k},E_{\nu,k}),E_{\nu,k})],$ (7)
@@
GANs are optimized using the so-called mini-max framework [45]. Hence, a successful model is a result of a balance between two losses: of the discriminator (5) and (6) and of the generator (7). $L_{D,1}$ in Eq. (5) is calculated on “true” samples obtained from NuWro, whereas the loss (6) is computed on “fake” data generated by G. The last loss (7) is evaluated on generated by model “fake” samples.
@@ II.4 Testing model’s quality
Figure 7: Top row: NuWro test sample (left) and GAN (right) generated samples for the QE scattering. Bottom row: Pull w/o tails per bin from NuWro and G-QE samples (left) and pull distribution (right).
@@
where $n_{\text{\sc NuWro}{},i}$ and $n_{gan,i}$ are the bin contents at the $i$-th bin, and $\sigma_{\text{\sc NuWro}{},i}^{2}=n_{\text{\sc NuWro}{},i}$ and $\sigma_{gan,i}^{2}=n_{gan,i}$ are statistical uncertainties. The bin contents of a histogram are expected to follow the Poisson distribution. As the G event distribution gets closer to NuWro’s, then Eq. (8) defines a random variable that follows a standard normal distribution $N(0,1)$ for large $n_{\text{\sc NuWro}{},i}$ and $n_{gan,i}$. For a random variable $x\sim N(0,1)$, the expected value of $|x|$ is
@@
The MAP metric is computed using histograms of the joint $E_{\mu}^{\prime}$ and $\theta^{\prime}$ distribution, as shown in Figs. 7 and 8. To reduce the bias coming from low statistics bins at distribution “tails”, we compute the MAP for bins for which $n_{nuw,i}>5$ and $n_{gan,i}>5$.
@@
In Fig. 9, we show how the EMD and MAP metrics change during the training of the QE and INC models. As can be noticed, after $100$ epochs, the optimization leads to a model for which the metrics begin to saturate. We stop the optimization using the MAP and EMD metrics. The best model minimizes both.
[…]
Figure 11: Event histograms, depending on $\theta^{\prime}$ (left column) and $\theta$ (right column) generated by the G-QE and NuWro test samples for neutrino energies =0.5absent0.5=0.5= 0.51111,  and $5$ GeV, shown in top, middle, and bottom rows, respectively.
@@ III Results
In this section, we compare the predictions made by the GAN model with the test data. To assess the quality of our models, we generated several datasets using NuWro. Each dataset was created for a specific value of neutrino energy, and we produced $1$ million events for each case. Additionally, we obtained test datasets (for QE and INC) similar to training, namely, one million events distributed uniformly in the energy range (300 MeV, 10 GeV).
In this section, we mainly discuss and compare GAN predictions with test datasets. In the appendix, we present the comparison between predictions of GAN and training datasets.
@@
Table 1: EMD and MAP (with and without tails) computed for the samples generated by the NuWro test sample and the G-QE for fixed neutrino energy $E_{\nu}$ and randomly sampled from a uniform $E_{\nu}$ distribution (“All” tag).
@@
Figure 12: Event histograms, depending on $p_{\mu,z}$ generated by the G-QE and NuWro test samples for neutrino energies 0.50.50.50.51111, and $3$ GeV, shown in top left, top right, and bottom figures, respectively.
The quality of the QE-GAN model and its dependence on neutrino energy is summarized in Table 1. We provide the values of the EMD and MAP metrics computed for several values of neutrino energies, between $0.5$ and $9$ GeV, as well as for a dataset with energies distributed uniformly in the energy region that covers the same energy domain as the training dataset. The performance of the G-QE model is excellent in all the considered cases.
In Fig. 7, we show the two-dimensional histograms and pull distribution of events generated by the NuWro and G-QE that are uniformly distributed in energies from $300$ MeV to $10$ GeV. In both GAN- and NuWro-generated samples, we clearly see the region where the events are concentrated. It corresponds to QE peak. The pulls per bin resemble statistical fluctuations, and their distribution approaches the Gaussian shape.
@@
Fig. 11 presents the dependence on $\theta^{\prime}$ and $\theta$ for neutrino energies of 0.50.50.50.51111, and $5$ GeV. Furthermore, we derive the longitudinal momentum as a composite variable (1) and show its event distribution in Fig. 12. The agreement between the GAN and the NuWro generated results is excellent. Note that for the comparison, we generated a new set of the NuWro events that were not used during training for all these tests.
Let us now discuss the INC model. In Fig. 8, we show the two-dimensional histograms and pull distribution of INC events generated by the NuWro and G-INC that are uniformly distributed in energies from $300$ MeV to $10$ GeV. Again, a good agreement is achieved between the NuWro and G-INC generator. Both GAN and NuWro reproduce the QE and $\Delta$ resonance peaks, see two accumulations of events in Fig. 8 (top panels). In Fig. 13, we plot the histograms of events generated by the G-INC and the NuWro MC generator for neutrino energies 11113333 and $5$ GeV for both $E^{\prime}_{\mu}$ and $E_{\mu}$.
Figure 13: Event histograms, depending on $E_{\mu}^{\prime}$ (left column) and $E_{\mu}$ (right column) generated by the G-INC and the NuWro test samples for neutrino energies $=1$, $3$, and $5$ GeV, shown in top, middle and bottom rows, respectively.
Figure 14: Event histograms, depending on $p_{\mu,z}$ (left column) and hadronic invariant mass $W$ (right column) generated by the G-INC and the NuWro test samples for neutrino energies $E_{\nu}=0.5$, $1$, and $5$ GeV, shown in top, middle and bottom rows, respectively.
To test if our GAN model can reproduce the QE and $\Delta(1232)$ resonance peaks equally well, let us now consider event distributions in hadronic invariant mass $W$ defined as
@@
where $\omega=E_{\nu}-E_{\mu}$ is the energy transferred to the nucleus, $M_{N}$ the average nucleon mass of the nucleon, and $Q^{2}$ the four-momentum transfer
@@
Figure 15: Two-dimensional inclusive distributions of NuWro events in $E_{\mu}^{\prime}$ and $\theta^{\prime}$ variables. Top: the training dataset for $E_{\nu}$ ranging from 950 MeV to 1050 MeV ($\sim 40$k events). Bottom: Testing dataset of one million events for $1$ GeV neutrino energy. Both figures show the approximate outline (based on histogram binning) of the $1$ GeV dataset distribution. The training dataset does not cover the whole kinematic range of the NuWro generator for neutrino energy around 1 GeV.
@@
It’s important to note that the test dataset covers a wider range of lepton kinematics than the training data. In Fig. 15, we display a histogram of the test data (INC) for neutrino energy at 1 GeV, alongside the training data for neutrino energies between $950$ MeV and $1050$ MeV. This comparison reveals that there are areas of lepton kinematics that are not represented in the training dataset but still are effectively reproduced by the GAN.